# Strong Screening Rules for Group-based SLOPE Models

### v1

@@ Abstract
Tuning the regularization parameter in penalized regression models is an expensive task, requiring multiple models to be fit along a path of parameters. Strong screening rules drastically reduce computational costs by lowering the dimensionality of the input prior to fitting. We develop strong screening rules for group-based Sorted L-One Penalized Estimation (SLOPE) models: Group SLOPE and Sparse-group SLOPE. The developed rules are applicable for the wider family of group-based OWL models, including OSCAR. Our experiments on both synthetic and real data show that the screening rules significantly accelerate the fitting process. The screening rules make it accessible for group SLOPE and sparse-group SLOPE to be applied to high-dimensional datasets, particularly those encountered in genetics.
@@ 1 Introduction
As the amount of data collected increases, the emergence of high-dimensional data, where the number of features ($p$) is much larger than the number of observations ($n$), is becoming increasingly common in fields ranging from genetics to finance. Performing regression and discovering relevant features on these datasets is a challenging task, as classical statistical methods tend to break down. The most popular approach to meeting this challenge is the Lasso [30], which has given rise to the general penalized regression framework
$\hat{\beta}(\lambda)=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{f(\beta)+\lambda J(\beta;v)\right\},$ (1)
@@
where $f$ is the loss function, $J$ is a convex penalty norm, $v\succeq 0$ are penalty parameters, and $\lambda>0$ is the regularization parameter.
@@
A key aspect of fitting a penalized model is to tune the value of the regularization parameter, which determines the level of sparsity in the fitted model. Several approaches exist for tuning this parameter, including cross-validation [8, 15] and exact solution path algorithms [10], but can often be computationally expensive. Generally, the model is fit along an $l$-length path $\lambda_{1}\geq\ldots\geq\lambda_{l}\geq 0$. Finding approaches for speeding up the model fitting process can lead to large computational savings when tuning. Screening rules are popular methods that reduce the dimensions of the input dimensionality by discarding irrelevant features before the fitting process. Discarding features, especially in high-dimensional settings, has a tremendous impact on the computational costs.
@@
Denoting the active set of non-zero parameters at $\lambda_{k+1}$ by $\mathcal{A}_{v}(\lambda_{k+1})=\{i\in\{1,\dots,p\}:\hat{\beta}_{i}(\lambda_{k+1})\neq 0\}$, the goal of a screening rule is to use the solution at $\lambda_{k}$ to recover a screened set of features, $\mathcal{S}_{v}(\lambda_{k+1})$, which is a superset of $\mathcal{A}_{v}(\lambda_{k+1})$. The screened set can then be used as input for calculating the fitted values. By finding the smallest possible screened set, that still contains the active set, the number of irrelevant features included in the optimization process is reduced.
There are two types of screening rules: safe and heuristic. Safe rules, as suggested by the name, provide guarantees that any variables discarded are in fact inactive at the solution. The seminal work of El Ghaoui et al. [11] introduced the Safe Feature Elimination Rule (SAFE) for the lasso, in which safe regions were constructed to eliminate non-active features. Other examples of safe rules are given in [2, 25, 34], including sample screening [28]. Heuristic rules, on the other hand, tend to discard more variables, but this increased screening efficiency can lead to active variables being incorrectly discarded. Examples of heuristic screening rules are given in [1, 13, 31]. Amongst the proposed heuristic rules, of particular interest is the strong screening rule framework developed for the lasso [31]. As strong screening rules can lead to violations, Tibshirani et al. [31] complemented their rules with a check of the Karush–Kuhn–Tucker (KKT) [16] optimality conditions to ensure no violations occur. There are also hybrid screening regimes, which combine safe and heuristic rules [32, 37].
@@
A strong screening rule is formulated through the KKT conditions for Equation 1, given by:
@@
If the gradient $\nabla f(\hat{\beta}(\lambda_{k+1}))$ was available, the set $\mathcal{A}_{v}(\lambda_{k+1})$ could be identified by checking the subdifferential of the norm at zero. That is, identifying for which variables
$-\nabla f(\beta)\in\lambda\partial J(\mathbf{0};v)=\{x\in\mathbb{R}^{p}:J^{*}(x;\lambda v)\leq 1\},$ (3)
where $J^{*}$ refers to the dual norm of $J$. The subdifferential at zero, $\partial J(\mathbf{0};v)$, can also be seen to be the unit ball of its dual norm. As the gradient at $\lambda_{k+1}$ is not available in practice, an approximation of the gradient is used to find a screened subset of the features, $\mathcal{S}_{v}(\lambda_{k+1})$. The screened variables are combined with the previously active ones to form the reduced input space, $\mathcal{E}_{v}(\lambda_{k+1})=\mathcal{S}_{v}(\lambda_{k+1})\cup\mathcal{A}_{v}(\lambda_{k})$. The fitted values are then calculated by solving Equation 1 using only the variables in $\mathcal{E}_{v}(\lambda_{k+1})$. Optimality of the fitted values is checked for using the KKT conditions (Equation 2), with violations added into $\mathcal{E}_{v}$. This ensures features are not discarded that should be active.
@@ 1.1 Contributions
Over the years, several adaptive extensions to the lasso have been proposed, including the Sorted L-One Penalized Estimation (SLOPE) method [4]. SLOPE applies the sorted $\ell_{1}$ norm $J_{\text{slope}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, where $v_{1}\geq\ldots\geq v_{p}\geq 0,|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$. One key advantage of SLOPE is its ability to control the variable false discovery rate (FDR) under orthogonal data. Both safe and strong rules have been proposed for SLOPE [3, 12, 17], as well as exact solution path algorithms [9, 24]. As SLOPE is a non-separable penalty, safe screening rules require repeated screening during the fitting process [17]. This can be a computational bottleneck, as the calculation of the safe regions required for screening are often expensive.
Strong screening rules, on the other hand, are usually more time efficient as the extra cost of the screening and KKT checks tend to be outweighed by the significant savings of fitting on a reduced input space [31]. Motivated by this, we introduce strong screening rules for two group-based extensions of SLOPE: Group SLOPE (gSLOPE) [6] and Sparse-group SLOPE (SGS) [14].
Group SLOPE (gSLOPE) is an extension of SLOPE for selecting groups, defined by the norm
such that $\beta^{(g)}\in\mathbb{R}^{p_{g}}$ is a vector of the group coefficients and $p_{g}$ denotes the group sizes [6]. The norm has ordered penalty weights $w_{1}\geq\ldots\geq w_{g}\geq 0$ (described in Appendix A.1) which are matched to $\sqrt{p_{1}}\|\beta^{(1)}\|_{2}\geq\ldots\geq\sqrt{p_{m}}\|\beta^{(m)}\|_{2}$.
Sparse-group SLOPE (SGS) was proposed as an extension to SLOPE and gSLOPE, applying both variable and group penalization for bi-level selection [14]. For $\alpha\in[0,1]$ with variable weights $v$ and group weights $w$ (described in Appendix B.1), the SGS norm is defined as
Both of these approaches were developed for the analysis of features that naturally belong to groups. This situation is frequently encountered in the analysis of genetics data, where genes are grouped into pathways for the completion of a specific biological task. The gSLOPE approach selects groups of features, whereas SGS performs bi-level selection. Both approaches have been shown to inherit SLOPE’s ability to control FDR under orthogonal data: gSLOPE controls the group FDR [6] and SGS controls both variable and group FDRs [14].
Section 2 introduces a sparse-group screening rule framework. Based on this, strong screening rules for gSLOPE and SGS are developed and presented in Sections 3 and 4. The proofs of the theorems and propositions presented in Sections 2 and 3 are provided in the Appendices A.2 and B.3. Through a series of simulations and the analysis of real datasets (Sections 5), the proposed screening rules are shown to significantly reduce computational runtime. Additionally, due to the reduced input dimensionality offered by screening rules, issues with convergence in the analysis of large datasets are shown to be alleviated. These improvements are achieved without affecting the solution optimality.
@@
The screening rules proposed in this manuscript are further applicable to group-based Ordered Weighted $\ell_{1}$ (OWL) models. SLOPE is a particular case of an OWL model through the definition of its weights. As the proposed screening rules only require that the penalty sequences ($v,w$) are ordered, they can be used for any group-based OWL models; for example, the Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) model [5]. Results for group-based OSCAR models are provided in Appendix E.
Notation.
@@
The feature space is divided into a set of $m$ non-overlapping groups, where $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$ denote the set of variable indices for each group. The set of active groups is denoted by $\mathcal{A}_{g}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}\neq 0\}$ whereas $\mathcal{Z}=\{i\in\{1,\dots,m\}:\|\hat{\beta}^{(i)}\|_{2}=0\}$ presents the inactive groups. The set of variable indices of the active and inactive groups are denoted by $\mathcal{G}_{\mathcal{A}}$ and $\mathcal{G}_{\mathcal{Z}}$, respectively. The cardinality of a vector $x$ is denoted by $\operatorname{card}(x)$. The operators $(\cdot)_{\downarrow}$ and $(\cdot)_{|\downarrow|}$ sort a vector into decreasing and decreasing absolute form. We use $\preceq$ to denote element-wise less than or equal to. The operator $\mathcal{O}(\cdot)$ returns the index of a vector sorted in decreasing absolute form. The cumulative summation operator applied on a vector is denoted by $\text{cumsum}(x)=[x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{\operatorname{card}(x)}x_{i}]$.
2 Screening rule framework
@@
  Input: $\lambda\in\mathbb{R}^{l}$, $\mathbf{X}\in\mathbb{R}^{n\times p},y\in\mathbb{R}^{n}$
@@
  Output: $\hat{\beta}(\lambda_{1}),\ldots,\hat{\beta}(\lambda_{l})\in\mathbb{R}^{p\times l}$
Sparse-group models apply both variable and group penalization, such that Equation 1 is extended to include an additional convex penalty function. Examples of sparse-group models include the Sparse-group Lasso (SGL) [29] and SGS [14]. Safe screening rules are available for SGL which perform bi-level screening, i.e. both group and variable screening [21, 33], but there are no strong rules which perform bi-level screening. The strong screening rule framework proposed by Tibshirani et al. [31] does not fully extend to sparse-group or non-separable penalties, whilst the strong screening rule derived for SGL in Liang et al. [19] applies only group-level screening.
Framework.
We introduce a new framework (Algorithm 1) for applying strong screening rules to sparse-group penalties that allow for bi-level screening, based on the framework introduced by Tibshirani et al. [31]. In the sparse-group framework, a screened set of groups is computed, $\mathcal{S}_{g}$, from the full input. An additional layer of screening is then performed to compute $\mathcal{S}_{v}$ using $\mathcal{S}_{g}$. This forms the reduced input set for fitting, $\mathcal{E}_{v}$. The aim of finding $\mathcal{E}_{v}$ is the same as for the lasso and SLOPE strong rules, but the framework first discards irrelevant groups. KKT checks are performed on $\mathcal{E}_{v}$ to ensure no violations have occurred. The framework applied to gSLOPE and SGS is presented in Appendix D.
Based on the framework described in Algorithm 1, the screening rules for SGS are derived (Section 4). The screening rule for gSLOPE (Section 3) follows a similar framework, where $\mathcal{S}_{v}$ is taken as all variables within the groups of $\mathcal{S}_{g}$. The KKT checks are then performed only on the groups.
By applying bi-level screening for SGS, a substantially larger proportion of variables are discarded, than if just group screening were applied (Figure 1). Using only group screening for SGS, without the additional variable screening step, resulted in the number of variables in the screened set scaling poorly with increased dimensionality. This demonstrates that bi-level screening effectively manages the scaling of SGS under rising dimensions.
(a) Applied to synthetic data (Section 5.1). Data generated under a linear model for $p=500,5000$. The results are averaged over $100$ repetitions.
(b) Applied to two real genetics datasets (Section 5.2), with groups formed using pathways. The results are averaged over the nine pathway collections.
Figure 1: The proportion of variables in $\mathcal{S}_{v}$ relative to the full input for SGS, shown for group and bi-level screening, plotted as a function of the regularization path, with $95\%$ confidence intervals.

### v2

@@ Abstract
Tuning the regularization parameter in penalized regression models is an expensive task, requiring multiple models to be fit along a path of parameters. Strong screening rules drastically reduce computational costs by lowering the dimensionality of the input prior to fitting. We develop strong screening rules for group-based Sorted L-One Penalized Estimation (SLOPE) models: Group SLOPE and Sparse-group SLOPE. The developed rules are applicable to the wider family of group-based OWL models, including OSCAR. Our experiments on both synthetic and real data show that the screening rules significantly accelerate the fitting process. The screening rules make it accessible for group SLOPE and sparse-group SLOPE to be applied to high-dimensional datasets, particularly those encountered in genetics.
@@ 1 INTRODUCTION
As the amount of data collected increases, the emergence of high-dimensional data, where the number of features ($p$) is much larger than the number of observations ($n$), is becoming increasingly common in fields ranging from genetics to finance. Performing regression and discovering relevant features on these datasets is a challenging task, as classical statistical methods tend to break down. The most popular approach to meeting this challenge is the lasso (Tibshirani, 1996), which has given rise to the general penalized regression framework
$\hat{\beta}(\lambda)\in\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{f(\beta)+\lambda J(\beta;v)\right\},$ (1)
@@
where $f$ is a differentiable and convex loss function, $J$ is a convex penalty norm, $v$ are penalty weights, and $\lambda>0$ is the regularization parameter.
@@
A key aspect of fitting a penalized model is to tune the value of $\lambda$ along an $l$-length path $\lambda_{1}\geq\ldots\geq\lambda_{l}\geq 0$. Several approaches exist for tuning this parameter, including cross-validation (Homrighausen and McDonald, 2018; Chetverikov et al., 2021) and exact solution path algorithms (Efron et al., 2004), but these can be computationally expensive. Screening rules help alleviate these costs by discarding variables that are inactive at the optimal solution, thus reducing input dimensionality prior to optimization.
@@
Denote the active set of coefficients at path point $\lambda_{k+1}$, for $k\in[l-1]:=\{1,\ldots,l-1\}$, by $\mathcal{A}_{v}(\lambda_{k+1})=\{i\in[p]:\hat{\beta}_{i}(\lambda_{k+1})\neq 0\}$. The goal of a (sequential) screening rule is to use the solution at $\lambda_{k}$ to recover a screened set of features, $\mathcal{S}_{v}(\lambda_{k+1})$, which is a superset of $\mathcal{A}_{v}(\lambda_{k+1})$. The screened set is then used as input for calculating the fitted values, leading to significant computational savings.
There are two types of screening rules: safe and heuristic. Safe rules are guaranteed to only discard inactive variables and mostly follow the Safe Feature Elimination (SAFE) framework (El Ghaoui et al., 2010), in which safe regions for variables are constructed. Other notable examples include Slores (Wang et al., 2014), the dome test (Xiang and Ramadge, 2012), and Dual Polytope Projections (DPP) (Wang et al., 2013), as well as sample screening (Shibagaki et al., 2016) and other examples given in Ogawa et al. (2013); Fercoq et al. (2015); Atamturk and Gomez (2020).
Heuristic rules tend to follow the strong screening rule framework, proposed by Tibshirani et al. (2010), and discard considerably more variables than safe rules. However, they can incorrectly discard active variables, so are complemented by a check of the Karush–Kuhn–Tucker (KKT) (Kuhn and Tucker, 1950) stationarity conditions. Other strong rules include Blitz (Johnson and Guestrin, 2015), SIS (Fan and Lv, 2008), and ExSIS (Ahmed and Bajwa, 2019). Hybrid schemes exist, using both safe and heuristic rules (Zeng et al., 2021; Wang and Breheny, 2022).
@@
A strong screening rule is formulated through the KKT stationarity conditions for Equation 1, given by
@@
If the gradient were available, the active set could be identified exactly by checking the subdifferential of the norm at zero:
$\nabla f(\beta)\in\lambda\partial J(\mathbf{0};v)=\{x\in\mathbb{R}^{p}:J^{*}(x;\lambda v)\leq 1\},$ (3)
where $J^{*}$ is the dual norm of $J$ and $\partial J(\mathbf{0};v)$ is the unit ball of the dual norm. So, $J^{*}(\nabla f(\beta);\lambda v)\leq 1$ indicates that $\beta=0$. As the gradient at $\lambda_{k+1}$ is not available, a model-specific approximation of the gradient is derived to find a screened subset of the features, $\mathcal{S}_{v}(\lambda_{k+1})$, such that $\mathcal{A}_{v}(\lambda_{k+1})\subset\mathcal{S}_{v}(\lambda_{k+1})$.
1.1 Screening Approaches for SLOPE
As the lasso is inconsistent under certain scenarios (Zou, 2006), several adaptive extensions have been proposed, including the Sorted L-One Penalized Estimation (SLOPE) model (Bogdan et al., 2015). SLOPE applies the sorted $\ell_{1}$ norm $J_{\text{slope}}(\beta;v)=\sum_{i=1}^{p}v_{i}|\beta|_{(i)}$, where $v_{1}\geq\ldots\geq v_{p}\geq 0,|\beta|_{(1)}\geq\ldots\geq|\beta|_{(p)}$. One key advantage of SLOPE is its ability to control the variable false discovery rate (FDR) under orthogonal data. Additional powerful properties include: it clusters strongly correlated features, it finds the minimum total squared error loss across different sparsity levels, removing the need for prior knowledge of sparsity, and it is asymptotically minimax (Figueiredo and Nowak, 2014; Su and Candès, 2016). All of these useful properties have meant that SLOPE has found widespread use in machine learning and genetics (Gossmann et al., 2015; Virouleau et al., 2017; Kremer et al., 2020; Frommlet et al., 2022; Riccobello et al., 2023).
Both safe (Bao et al., 2020; Elvira and Herzet, 2021) and strong (Larsson et al., 2020) rules have been proposed for SLOPE, as well as exact solution path algorithms (Nomura, 2020; Dupuis and Tardivel, 2023). As SLOPE is a non-separable penalty, safe screening requires repeated screening during optimization, which is expensive due to the repeated dual norm evaluations required for the safe regions (Larsson et al., 2020).
Group-based SLOPE Models
In genetics, the analysis of grouped features is frequently encountered, as genes are grouped into pathways for the completion of a specific biological task. To use this grouping information, SLOPE has been extended to group (gSLOPE) and sparse-group (SGS) regression.
For a set of $m$ non-overlapping groups, $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$ of sizes $p_{1},\ldots,p_{m}$, Group SLOPE (gSLOPE) (Gossmann et al., 2015; Brzyski et al., 2019) is given by
such that $\beta^{(g)}\in\mathbb{R}^{p_{g}}$ is a vector of the group coefficients. The norm has ordered penalty weights $w_{1}\geq\ldots\geq w_{g}\geq 0$ (described in Appendix A.1) which are matched to $\sqrt{p_{1}}\|\beta^{(1)}\|_{2}\geq\ldots\geq\sqrt{p_{m}}\|\beta^{(m)}\|_{2}$.
Sparse-group SLOPE (SGS) (Feser and Evangelou, 2023) was further proposed as a convex combination of SLOPE and gSLOPE for concurrent variable and group selection. For $\alpha\in[0,1]$, with weights $(v,w)$ (described in Appendix B.1), the norm is given by
Both approaches control the FDR under orthogonal data: gSLOPE at the group-level (Brzyski et al., 2019) and SGS at both levels (Feser and Evangelou, 2023). SGS has been found to outperform other methods at selection and prediction (Feser and Evangelou, 2023).
@@ 1.2 Contributions
No screening rules exist for group-based SLOPE models. The strong screening rule framework introduced by Tibshirani et al. (2010) facilitated the extension from the lasso to the group lasso by requiring only knowledge of the dual norm. However, this framework is restricted to separable penalties, making it unsuitable for SLOPE. Instead, in Larsson et al. (2020), the subdifferential of SLOPE is used to derive strong rules.
Motivated by this, we propose a new strong screening framework for sparse-group norms (Section 2), used to develop screening for gSLOPE (Section 3) and SGS (Section 4) (with proofs of the results provided in Appendices A.3 and B.3). The framework applies two screening layers, drastically reducing dimensionality (Figure 1). The screening requires knowledge of the subdifferentials, necessitating a general derivation for gSLOPE (Theorem 3.1).
The choice of strong screening over safe is motivated by two main reasons. First, strong rules discard significantly more variables than safe rules (Tibshirani et al., 2010). Second, safe rules require calculating the dual norm set. For SGS, this is a sum of convex sets. Determining if a point lies in this set requires knowledge of the summation’s decomposability, which is a difficult task (Wang and Ye, 2014). This challenge is addressed in Section 4.1.
The analysis of synthetic and real datasets shows that our proposed screening rules considerably improve runtime (Section 6). The reduced input dimensionality from the screening also eases convergence issues with large datasets. These improvements are achieved without affecting solution optimality.
Figure 1: The proportion of variables in $\mathcal{S}_{v}$ relative to $p$ for group-only and bi-level screening applied to SGS, plotted along the regularization path with 95% confidence intervals. Synthetic data was generated under a linear model for $p=500,5000$ (Section 6.1), with results averaged over 100 repetitions.
@@
As the proposed screening rules only require that the penalty sequences ($v,w$) are ordered, they apply to the wider class of group-based Ordered Weighted $\ell_{1}$ (OWL) models. Popular special cases of OWL models include SLOPE and the Octagonal Shrinkage and Clustering Algorithm for Regression (OSCAR) model (Bondell and Reich, 2008) (the description of group-based OSCAR models is provided in Section 5).
Notation
@@
The sets of active and inactive groups are given by $\mathcal{A}_{g}=\{g\in[m]:\|\hat{\beta}^{(g)}\|_{2}\neq 0\}$ and $\mathcal{Z}=\{g\in[m]:\|\hat{\beta}^{(g)}\|_{2}=0\}$. Their corresponding set of variable indices are denoted by $\mathcal{G}_{\mathcal{A}}$ and $\mathcal{G}_{\mathcal{Z}}$. The cardinality of a vector $x$ is denoted by $\operatorname{card}(x)$. The operators $(\cdot)_{\downarrow}$ and $(\cdot)_{|\downarrow|}$ sort a vector into decreasing and decreasing absolute form. We use $\preceq$ to denote element-wise inequality signs. The operator $\mathcal{O}(\cdot)$ returns the index of a vector sorted into decreasing absolute form. The cumulative summation operator applied on a vector is denoted by $\text{cumsum}(x)=[x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{\operatorname{card}(x)}x_{i}]$.
2 SPARSE-GROUP STRONG SCREENING
Sparse-group models, such as SGS and the sparse-group lasso (SGL) (Simon et al., 2013), apply both variable and group penalization so that bi-level screening is possible. Safe rules that perform bi-level screening exist for SGL (Wang and Ye, 2014; Ndiaye et al., 2016a), but there are no such strong rules. The strong screening framework (Tibshirani et al., 2010) does not extend to sparse-group or non-separable norms, and the strong rule derived for SGL in Liang et al. (2022) applies only group-level screening.
Framework
We introduce a new sparse-group framework (Algorithm 1), based on the strong framework by Tibshirani et al. (2010), for applying strong screening to sparse-group norms, allowing for bi-level screening. By applying bi-level screening for SGS, a substantially larger proportion of variables are discarded than with group screening alone (Figure 1).
First, a screened set of groups is computed, $\mathcal{S}_{g}$. An additional layer of screening is then performed to compute $\mathcal{S}_{v}$ using $\mathcal{S}_{g}$. This, combined with the previously active variables, forms the reduced input set for fitting, $\mathcal{E}_{v}$. KKT checks are performed on $\mathcal{E}_{v}$ to ensure no violations have occurred (Appendices A.4 and B.4). Based on this framework, the SGS rules are derived in Section 4. The screening for gSLOPE (Section 3) also uses this framework but does not perform the variable screening and instead takes $\mathcal{S}_{v}$ as all variables in the groups of $\mathcal{S}_{g}$. The KKT checks are then performed only on the groups. Appendix D describes the implementation of the framework for gSLOPE and SGS.
The main cost of the framework is calculating the fitted values. For $t$ iterations of ATOS (a proximal algorithm used to fit SGS, see Section 6.1), the convergence rate is $O(1/t)$ (Pedregosa and Gidel, 2018) and we expect a time complexity of $O(tp^{2})$ for a proximal algorithm (Zhao and Huo, 2023).
@@ 2 SPARSE-GROUP STRONG SCREENING
  Input: $(\lambda_{1},\ldots,\lambda_{l})\in\mathbb{R}^{l}$, $\mathbf{X}\in\mathbb{R}^{n\times p},y\in\mathbb{R}^{n}$
@@
  Output: $\hat{\beta}(\lambda_{1}),\ldots,\hat{\beta}(\lambda_{l})\in\mathbb{R}^{p}$